Circular No. 6423

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996ai IN NGC 5005 A. V. Filippenko, University of California (UC) at Berkeley, reports:

“Preliminary inspection of an uncalibrated, somewhat noisy CCD spectrum (range 360–760 nm), obtained by N. P. Vogt (UC, Santa Cruz) and M. A. Bershady (Pennsylvania State University) on June 20 UT with the Lick Observatory 3-m Shane reflector, suggests that SN 1996ai (IAUC 6422) is a supernova, probably of type Ia prior to maximum brightness. There is a broad absorption trough at 614 nm, the defining feature of type-Ia supernovae, but this might instead be redshifted Hα absorption in a type-II object (although the corresponding Hα emission line is weak or absent). No other deep features are visible; weak ones are consistent with the type-Ia classification. The object appears to be quite red, perhaps due to extinction; a relatively strong (EW = 0.3 ± 0.1 nm), narrow Na D absorption line at the redshift of the host galaxy is visible. Crude CCD magnitude estimates from H. Teplitz (UC, Los Angeles) on June 20 with the Lick 1-m Nickel telescope: B = 16.5, V = 15.0, R = 14.5, I = 13.5.”

CM DRACONIS E. Guinan, G. McCook, and S. Wright, Villanova University, write:

“We have made a possible photometric detection of a planetary transit eclipse for the eclipsing-binary star CM Dra — selected as a planet-search target because its orbital plane is seen almost exactly edge-on (Doyle et al. 1995, BAAS 27, 1382). Photoelectric photometry of CM Dra has been conducted from Mt. Hopkins since 1995 using the Four College Consortium 0.8-m APT. The binary (P = 1.268 days) has been well behaved, showing the expected mutual eclipses of the two stellar components and nearly constant light between the narrow eclipse phases. During a 3.5-hr observation beginning June 1.177 UT, CM Dra was fainter by 0.08 mag in the I band; if we assume that this I decrease is due to a planet transiting one of the two dM5 stars, it would have a diameter of ∼ 0.85 that of Jupiter. From the lower limit to the transit duration, the possible planet would have an orbital period of at least several months. Though this decrease in light may be due to starspot development, photometry the day before and after the dimming event showed the system to be at its usual brightness. Evidence of the existence of a planet around CM Dra is being investigated by searching for small periodic deviations of the eclipse timings; a planet of Jupiter’s mass at 1 AU would produce ∼ 17-s periodic shift (observable with accurate photometry) in the arrival times of the eclipses of the stars.”

1996 June 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green